# Dynamical breaking of inversion symmetry, strong second harmonic generation, and ferroelectricity with nonlinear phonons

### v2

@@ Introduction
We note that phonons with frequencies close to $\Omega_{0}/2$ can be exploited to resonantly enhance the otherwise off-resonant driving. Consider an auxiliary, IR active phonon mode $P_{A}$, such that it couples to $Q_{x}$ via a term
$H_{PQ}=\lambda Q_{x}P_{A}.$
This coupling preserves the original inversion symmetry of the system and leads to $\lambda P_{A}$ taking over the role of the electric field in Eq. (3). $P_{A}$ can then be a regular IR active mode following (to zeroth order in $\lambda$) the equation of motion $\ddot{P}_{A}+2\gamma_{A}\dot{P}_{A}+\omega_{A}P_{A}=Z_{A}E_{x}\cos\left(\omega_{A}t\right)$, such that it accumulates the energy of the electric field over a number of $\sim\omega_{A}/\gamma_{A}$ cycles. Due to the inherent nonlinear blue-shift of the effective resonance frequency $\tilde{\Omega}_{0}$ [Eq. (11)], the driving power can be adjusted such that the resonance frequency $\omega_{A}$ of the auxiliary phonon mode $P_{A}$ exactly hits $\omega_{R}=\tilde{\Omega}_{0}/2$, very similar to the situation depicted in Fig. 1 a), where, the driving frequency is fixed to $0.6\Omega_{0}$, while the driving amplitude is increased. Around $E_{x}=0.4\,\text{\AA }\sqrt{u}\Omega_{0}^{2}/Z_{x}$, the resonance condition is fulfilled and the systems enters the symmetry breaking regime. Thus the auxiliary mode does not have to be located at exactly half the resonant frequency of the $Q_{x}$ mode, rather, the blue shift can be exploited to access the instability at the twice the frequency of the auxiliary mode by adjusting the driving strength.
Collective instability of the x and y modes
Having studied the symmetry breaking instability for a single phonon component driven by linearly polarized light, we now turn to the full chiral system consisting of modes $Q_{x}$ and $Q_{y}$ described by Eqs. (3), (4). Numerically, we observe that the instability intervals are larger if $E_{x}\neq E_{y}$, i.e. the driving electromagnetic field is eliptically polarized. To rationalize this observation, we perform a stability analysis for the two-component equations (3), (4) following Ref. (Landau_Lifshitz_Mechanics, ).
$\ddot{Q}_{x,\mathrm{even}}+2\gamma\dot{Q}_{x,\mathrm{even}}+\left[\tilde{\Omega}_{x}^{2}+\frac{\alpha}{2}\left(3F_{x,1}^{2}-F_{y,1}^{2}\right)\cos\left(2\omega t\right)\right]Q_{x,\mathrm{even}}+\alpha F_{x,1}F_{y,1}\sin\left(2\omega t\right)Q_{y,\mathrm{even}}$ $=0$
$\ddot{Q}_{y,\mathrm{even}}+2\gamma\dot{Q}_{y,\mathrm{even}}+\left[\tilde{\Omega}_{y}^{2}-\frac{\alpha}{2}\left(3F_{y,1}^{2}-F_{x,1}^{2}\right)\cos\left(2\omega t\right)\right]Q_{y,\mathrm{even}}+\alpha F_{x,1}F_{y,1}\sin\left(2\omega t\right)Q_{x,\mathrm{even}}$ $=0,$ (14)
with $\tilde{\Omega}_{x/y}^{2}=\Omega_{x/y}^{2}+\alpha\left(3F_{x/y}^{2}+F_{y/x}^{2}\right)/2$, where $F_{y}$ is defined analogously to $F_{x}$ in Eq. (8). As above, we expect parametric resonances near $2\omega=\tilde{\Omega}_{x/y}$, where $\tilde{\Omega}_{x}\neq\tilde{\Omega}_{y}$ for $F_{x}\neq F_{y}$. For now, let us choose the case $2\omega=\tilde{\Omega}_{x}$, such that the instability occurs for the $Q_{x}\left(t\right)$ component.
The oscillating terms in Eqs. (14) couple harmonics with frequencies $2\omega$, $4\omega$, … and DC terms. For the stability analysis, we therefore choose the ansatz
$Q_{i,\mathrm{even}}$ $=a_{i,1}\sin\left(2\omega t\right)+a_{i,2}\sin\left(4\omega t\right)$
$\quad+b_{i,0}+b_{i,1}\cos\left(2\omega t\right)+b_{i,2}\cos\left(4\omega t\right).$ (15)
Furthermore, we neglect $\gamma$ for the duration of this analysis. While $\gamma$ determines the instability threshold amplitudes of the electromagnetic fields [see Eq. (13)], its effects become less important for driving amplitudes above the threshold, i.e., for any driving amplitude, $\gamma$ can be always chosen small enough that our analysis is accurate. We again search for the instability window for the detuning $\Delta=2\omega-\tilde{\Omega}_{x}$, such that the mode amplitudes $a_{i,n}$ and $b_{i,n}$ grow exponentially for $\Delta_{\mathrm{min}}<\Delta<\Delta_{\mathrm{max}}$. At $\Delta=\Delta_{\mathrm{max}/\mathrm{min}}$, the amplitudes will be constant. The boundaries of the instability interval $\Delta_{\mathrm{max}/\mathrm{min}}$ are then found by inserting the ansatz (15) into Eqs. (14) and assuming that $a_{i,n}$ and $b_{i,n}$ are indeed constant (Landau_Lifshitz_Mechanics, ). After a lengthy calculation, in which we compare the coefficients of different harmonics after inserting the ansatz (15) into Eqs. (14), we find that, to fourth order in $F_{x,1}$ and $F_{y,1}$
$\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ $=F_{y,1}^{4}\left(\frac{\alpha^{2}}{16\Omega_{0}^{3}}+\frac{287\alpha^{4}F_{x,1}^{4}}{576\Omega_{0}^{7}}+\frac{47\alpha^{3}F_{x,1}^{2}}{192\Omega_{0}^{5}}\right)$
$\quad+F_{y,1}^{2}\left(\frac{77\alpha^{3}F_{x,1}^{4}}{192\Omega_{0}^{5}}-\frac{5\alpha^{2}F_{x,1}^{2}}{8\Omega_{0}^{3}}\right)+\frac{9\alpha^{2}F_{x,1}^{4}}{16\Omega_{0}^{3}}.$ (16)
The full result is too long to be quoted here but is easily found using computer algebra. Eq. (16) is valid for small $F_{x}$ and $F_{y}$, i.e. for small driving amplitudes. It is interesting to study the behavior of $\Delta$ close to $F_{y,1}=F_{x,1}$, i.e. for nearly perfect circular polarization. Writing $F_{y,1}=F_{x,1}+F_{\epsilon}$, we find
$\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ $=-F_{\epsilon}^{3}\left(\frac{9\Omega_{0}}{8F_{x,1}^{3}}+\frac{\alpha^{2}F_{x,1}}{\Omega_{0}^{3}}+\frac{51\alpha}{16F_{x,1}\Omega_{0}}\right).$ (17)
Notice that for $F_{y,1}\geq F_{x,1}$ (we choose both amplitudes positive w.l.o.g.), we have $\Delta_{\mathrm{max}}\leq\Delta_{\mathrm{min}}$, which indicates that the system is stable. For $F_{y,1}>F_{x,1}$, the $Q_{x}$ and $Q_{y}$ components switch places, and the instability occurs for $2\omega=\tilde{\Omega}_{y}$. The analysis for this case is completely analogous with $F_{x,1}$ and $F_{y,1}$ , as well as $\Omega_{x}$ and $\Omega_{y}$ interchanged. We therefore conclude that, in general, the instability occurs either for the $Q_{x}$ or the $Q_{y}$ component, depending on whether $F_{x,1}$ or $F_{y,1}$ is larger.
Figure 2: The instability window $\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ [see the discussion above Eq. (16)] as a function of $F_{y,1}/F_{x,1}$ is plotted for $F_{x,1}=0.1$. The full solution for the ansatz of Eq. (15) is plotted as a blue solid line in the main figure and in the inset. The dashed red line indicates the approximation of Eq. (16). The dashed green line in the inset shows the approximation of Eq. (17), which is valid close to $F_{y,1}=F_{x,1}$. Stable regions, where $\Delta_{\mathrm{max}}\leq\Delta_{\mathrm{min}}$ are marked yellow [see discussion below Eq. (17)]. We conclude that the system is stable at $F_{x,1}=F_{y,1}$ which holds for driving with perfectly circularly polarized light, i.e. $E_{x}=E_{y}$. Some amount of ellipticity of the driving electromagnetic field is necessary to access the instability.
The expansions of Eqs. (16) and (17) in comparison to the full result are shown in Fig. 2 for small amplitudes $F_{x}$, $F_{y}$. If driven with perfectly circularly polarized light with $E_{x}=E_{y}$, resulting in $F_{x}=F_{y}$, the time-dependent terms in Eqs. (14) cancel each other, and we find $\Delta_{\mathrm{min}}=\Delta_{\mathrm{max}}$; the symmetry breaking state cannot be reached. Thus, to induce the symmetry breaking transition, elliptically polarized light must be used  [^3] We note, that this must not necessarily be the case for other, less generic types of nonlinearities than the one considered here.
@@
In Fig. 1 d), we plot the numerical solutions for $Q_{x}\left(t\right)$, $Q_{y}\left(t\right)$ for elliptically polarized light with $E_{y}=0.25E_{x}$. The resulting Lissajous trajectory breaks the inversion symmetry of the Hamiltonian (1), due to the large second harmonic component of $Q_{x}\left(t\right)$.
@@
Beyond the instability at $\omega\approx\Omega_{0}/2$ that we have studies so far, higher order, inversion symmetry breaking instabilities at frequencies $\omega\approx\Omega_{0}/n$, where $n$ is an even number, can be induced. These instabilities generate higher order even harmonics. In general, the required threshold driving powers are larger for higher order parametric instabilities, and grow according to $E_{x,*}\sim\gamma^{1/2n}$ (Landau_Lifshitz_Mechanics, ). Fig. 1 e) shows the phonon trajectories for the $n=4$ instability, where $Q_{x}\left(t\right)$ exhibits a strong fourth harmonic component.
Non-degenerate chiral modes
@@
Finally, we investigate dynamical symmetry breaking for non-degenerate chiral Phonons. A toy-model with split frequencies for phonons of opposite chiralities is obtained by substituting $P_{i}\rightarrow P_{i}-\kappa A_{i}$ in the Hamiltonian of Eq. (1). Here $\mathbf{A}=B_{\mathrm{eff}}\left[-Q_{y},Q_{x},0\right]$ takes the role of an effective magnetic vector potential acting on the motion of the phonon components. To linear order in $\kappa$, the above substitution is equivalent to adding the term $\kappa\mathbf{B}_{\mathrm{eff}}\cdot\mathbf{L}$ to the Hamiltonian (1), i.e.
$H\rightarrow H+\kappa\mathbf{B}_{\mathrm{eff}}\cdot\mathbf{L},$ (18)
where $\mathbf{L}=\left(Q_{x}P_{y}-Q_{y}P_{x}\right)\hat{\mathbf{e}}_{z}$ is the phonon angular momentum and $\mathbf{B}_{\mathrm{eff}}=\left[0,0,B_{\mathrm{eff}}\right]$. Solving the linearized equations of motion, we find the phonon eigenfrequencies
@@
We find that the symmetry-breaking instability described above can also be achieved with non-degenerate chiral phonons. The right- and left-handed modes can be accessed separatly, depending on the polarization of the driving electromagnetic field. Because of the driving-induced blue-shift, the instabilty can be accessed for the two non-degenerate modes at the same frequency, but at different driving powers. In agreement with the results for degenerate chiral phonons discussed above [see Eq. (17)], we find that perfectly circular polarized light is ineffective in inducing the symmetry breaking. A certain amount of ellipticity is necessary. We present these results in Fig. 3.
Figure 3: Symmetry breaking with non-degenerate chiral phonons [see Eq. (18)]. The phonon frequencies are split according to Eq. (19): $\Omega_{+}$ corresponds to right-handed motion, while $\Omega_{-}$ corresponds to a left-handed rotation. The two modes are accessed with light of opposite polarizations, fitting their respective sense of motion. To excite the $\Omega_{+}$ mode, we use $\mathbf{E}=E_{0}\left[-\left(1-\delta\right)\cos\omega t,\sin\omega t,0\right]$, and for the $\Omega_{-}$ mode, $\mathbf{E}=E_{0}\left[\cos\omega t,\left(1-\delta\right)\sin\omega t,0\right]$, with $\delta=0.25$ and $\omega=0.62\Omega$. As for degenerate chiral phonons, a slight detuning from circularity $\delta$ is necessary, in order to trigger the instability at half the resonance frequency [see Eq. (17)]. The figure combines the results of two runs, in which the two chiralities were simulated separately. The resonance frequencies exhibit a driving amplitude dependent blue-shift, such that the instability occurs at different powers, for the two chiralities.
@@ Conclusion
In conclusion, we have described a new, symmetry breaking steady state for driven chiral, nonlinear phonons. This state is characterized by strong second harmonic generation and by the emergence of a ferroelectric response. These effects, being forbidden by the inversion symmetry of the underlying lattice, can serve as sharp experimental signatures of inversion symmetry breaking in the steady state and the effects presented in this manuscript. Beyond possible applications for second harmonic generation, rectification, and driven on-demand ferroelectricity, the study of interactions of chiral phonons in the newly described state with electrons and other collective modes (e.g. magnons (kahana2024_chiral_rectification, )) offers an intriguing avenue for uncovering novel out-of-equilibrium correlated states and dynamical phase transitions.

### v3

@@ Introduction
Auxiliary modes can be used to further simplify reaching the symmetry-breaking steady state. Infrared-active phonons with frequencies $\omega_{A}$ close to $\Omega_{0}/2$ can be exploited to resonantly enhance the otherwise off-resonant driving: consider a phonon mode $Q_{A}$ with a coupling term
$H_{qQ}=\lambda Q_{x}Q_{A}.$ (16)
This coupling preserves the original inversion symmetry of the system and leads to $\lambda Q_{A}$ taking over the role of the electric field in Eq. (A 7). The mode $q_{A}$ can be pumped resonantly, such that the energy of the drive is stored over $\sim\omega_{A}/\gamma_{A}$ cycles, resulting in a large amplitude oscillation of $Q_{A}$. Due to the nonlinear blue-shift of $\tilde{\Omega}_{0}$ [Eq. (10)], the frequency of the auxiliary phonon $\omega_{A}$ can be adjusted by choosing a suitable driving amplitude, such that it exactly hits $\omega_{A}=\tilde{\Omega}_{0}/2$.
Raman-active modes can be used in a similar way. Here the lowest order, inversion symmetric coupling Hamiltonian is given by
$H=AQ_{R}Q_{x}^{2}.$ (17)
$Q_{R}$ is an auxiliary Raman mode. As shown in Appendix Appendix.3, including the coupling of Eq. (17) again leads to an equation of motion for $Q_{x}$ that can be mapped onto the original Eq. (A 7). Thus Raman active modes can, too, be used to drive the system into the symmetry-breaking state. This might me advantageous, since driving these modes allows for greater optical penetration depths compared to infrared-active modes.
Nonlinear, circularly polarized phonons
Figure 3: The Lissajous trajectory of phonon coordinates $Q_{x}\left(t\right)$ and $Q_{y}\left(t\right)$ when driven into the symmetry-breaking state using elliptically polarized light with $\mathbf{E}=E_{x}\left[\cos\left(\omega t\right),0.25\sin\left(\omega t\right)\right]$. The inversion symmetry of Eq. (18) is broken dynamically.
We now investigate the effects of dynamical symmetry breaking on circularly polarized phonons. While sometimes such phonons are referred to as chiral (kahana2024_chiral_rectification, ), this terminology has been questioned (juraschek2025chiral_phonons_rev, ). We will discuss both, degenerate circularly polarized optical phonons (kahana2024_chiral_rectification, ; cheng2020_dirac_semi_phonon_zeeman, ; mustafa2025origin_phonon_zeeman_mos2, ), as well as phonons with split frequencies for right- and left-handed motion (zhang2015chiral_phonons_hexagonal, ; zhu2018observation_chiral_phonons, ; ishito2023truly_chiral_phonons, ; ueda2023chiral_phonons_xrays, ).
A simple model for nonlinear, degenerate, circularly polarized phonons is given by the Hamiltonian (kahana2024_chiral_rectification, )
$H=\frac{P_{x}^{2}}{2}+\frac{P_{y}^{2}}{2}+\frac{\Omega_{0}^{2}}{2}\left(Q_{x}^{2}+Q_{y}^{2}\right)+\frac{\beta}{4}\left(Q_{x}^{2}+Q_{y}^{2}\right)^{2}+V_{\mathrm{l-m}}.$ (18)
Here, $Q_{x}$ and $Q_{y}$ are the coordinates of two orthogonal phonon modes. The light-matter coupling $V_{\mathrm{l-m}}$ is given by
where $\mathbf{p}_{n}=\mathbf{Z}_{n}Q_{n}$ are the electric dipole moments of the phonon components, with effective charges $\mathbf{Z}_{n}$. For simplicity, we assume $\mathbf{Z}_{n}\propto\hat{\mathbf{e}}_{n}$. The equations of motion for elliptically polarized light are given in AppendixAppendix.4, where we added damping at rate $\gamma$.
Solving the equations of motion (A 7) numerically, we observe that the instability intervals are larger for elliptical polarization with $E_{x}\neq E_{y}$. A stability analysis for the system following Ref. (Landau_Lifshitz_Mechanics, ), for which we refer to AppendixAppendix.4, rationalizes this observation.
@@
The result of a numerical simulation of the phonon trajectory in the symmetry-breaking state is shown in in Fig. 3 for driving with elliptically polarized light where $E_{y}=0.25E_{x}$. The resulting Lissajous trajectory breaks the inversion symmetry of the Hamiltonian (18), due to the large second harmonic component of $Q_{x}\left(t\right)$.
@@
We note that beyond the instability at $\omega\approx\Omega_{0}/2$, higher order, inversion symmetry breaking instabilities at frequencies $\omega\approx\Omega_{0}/n$, where $n$ is an even number can be induced. These generate more complex Lissajous-figures, however the required threshold driving amplitudes grow according to $E_{x,*}\sim\gamma^{1/2n}$ (Landau_Lifshitz_Mechanics, ).
@@
Finally, we investigate dynamical symmetry breaking for non-degenerate circularly polarized phonons (zhang2015chiral_phonons_hexagonal, ) (also called “false chiral phonons” (juraschek2025chiral_phonons_rev, )). Such phonons can appear in systems with broken time-reversal symmetry (bonini2023CrI3_chiral_phonons, ) and have been observed in monolayer TMDs (zhu2018observation_chiral_phonons, ). A toy-model with split frequencies for phonons of opposite chiralities, is obtained by substituting $P_{i}\rightarrow P_{i}-\kappa A_{i}$ in the Hamiltonian of Eq. (18) (chen2019chiral_chen_review, ). Here $\mathbf{A}=B_{\mathrm{eff}}\left[-Q_{y},Q_{x},0\right]$ takes the role of an effective magnetic vector potential. To linear order in $\kappa$, the above substitution is equivalent to adding the term $\kappa\mathbf{B}_{\mathrm{eff}}\cdot\mathbf{L}$ to the Hamiltonian (18), where $\mathbf{L}=\left(Q_{x}P_{y}-Q_{y}P_{x}\right)\hat{\mathbf{e}}_{z}$ is the phonon angular momentum and $\mathbf{B}_{\mathrm{eff}}=\left[0,0,B_{\mathrm{eff}}\right]$. Solving the linearized equations of motion, we find the phonon eigenfrequencies
@@
We find that the symmetry-breaking instability described above can also be achieved with non-degenerate chiral phonons and present these results in Fig. 4.
Figure 4: Symmetry breaking with non-degenerate chiral phonons [see Eq. (20)]. The phonon frequencies are split according to Eq. (20): $\Omega_{\pm}$ corresponds to right/left-handed motion. The two modes are accessed with light of opposite polarizations. To excite the $\Omega_{+}$ mode, we use $\mathbf{E}=E_{0}\left[-\left(1-\delta\right)\cos\omega t,\sin\omega t,0\right]$, and for the $\Omega_{-}$ mode, $\mathbf{E}=E_{0}\left[\cos\omega t,\left(1-\delta\right)\sin\omega t,0\right]$, with $\delta=0.25$ and $\omega=0.62\Omega$. As for degenerate chiral phonons, a slight detuning from circularity $\delta$ is necessary, in order to trigger the instability at half the resonance frequency [see Eq. (A 11)]. The figure combines the results of two runs, in which the two chiralities were simulated separately. The resonance frequencies exhibit a driving amplitude dependent blue-shift, such that the instability occurs at different powers, for the two chiralities.
@@ Conclusion
We have described a new, inversion-symmetry-breaking steady state for driven nonlinear phonons. This state is characterized by strong second harmonic generation and by the emergence of ferroelectric behavior. These responses – forbidden by the inversion symmetry of the underlying lattice – can serve as sharp experimental signatures of the physics presented in this manuscript.
Ferroelectrics above the transition (ginzburg1949theory_ferroelectrics, ; cochran1959_cochran_ferroelectrics1, ; cochran1960_cochran_ferroelectrics2, ) are natural platforms for realizing the effects. The softening phonon modes responsible for the transition will have low frequencies and strong anharmonicity (scott1974soft_modes_review, ; pal2021origin_thz_soft_modes, ; alyabyeva2021_barium_hexaferrite_soft_modes, ), resulting in low threshold powers $E_{x,*}$ as shown in Eq. (12).
The symmetry-breaking state can be induced with flat-top pulses as short as five optical cycles, and at reachable field strengths (Table1). Beyond possible applications for second harmonic generation, rectification, and driven, on-demand ferroelectricity, the study of interactions of phonons in the newly described state with electrons and collective modes offers a route to novel out-of-equilibrium states.
@@ Acknowledgements.
Appendix
Appendix.1 Achieving symmetry breaking with pulsed excitations
We perform simulations to show the symmetry-breaking steady state can be reached with short pulsed signals. We confine the oscillating electric field in Eq. (3) to a window of length $T_{p}=5\cdot 2\pi/\omega$. This mimics a flat-top pulse (see Fig.A 1) whose amplitude is constant throughout $T_{p}$. The constant amplitude is necessary, because due to the blue-shift of $\tilde{\Omega}_{0}\left(E_{x}\right)$ both amplitude and frequency of the drive have to be matched with each other (see Fig. 1). Fig. 1 shows the flat-top pulse and the response of the phonon mode $Q_{x}\left(t\right)$.
Figure A 1: Reaching the symmetry-breaking state is possible with a short flat-top pulse of only five optical cycles. The upper plot shows the driving electric field, while the lower plot shows the response with visible second harmonic and DC offset components. When the pulse ends, the system switches to unforced oscillations at frequency $\Omega_{0}$. The of $Q_{x}\left(t\right)$ is show in Fig.1 of the main text.
Appendix.1.1 A note on initial conditions and symmetry breaking
We started our simulation runs at $t=0$ and used the forcing term $E_{x}\left(t\right)=E_{0,x}\cos\left(\omega t\right)$, effectively this corresponds to $E_{x}\left(t\right)\sim\Theta\left(t\right)\cos\left(\omega t\right)$. The response of Eq. (3) to this forcing necessarily consists of particular solution oscillating at $\omega$ and persisting for $t\rightarrow\infty$, and a homogeneous solution oscillating at $\tilde{\Omega}_{0}$ that fixes the boundary conditions at $t=0$, and decays due to damping. The homogeneous solution has both finite $Q_{\mathrm{even}}$ and $Q_{\mathrm{odd}}$ components since $\tilde{\Omega}_{0}$ is not necessarily a multiple of $\omega$. It thus initializes the parametric instability in Eq. (9), which gives $Q_{\mathrm{even}}\left(t\right)=0$ for $Q_{\mathrm{even}}\left(0\right)=0$ and $\dot{Q}_{\mathrm{even}}\left(0\right)=0$, acting as an initial forcing for $Q_{\mathrm{even}}$, and determining the sign of the symmetry breaking.
Appendix.2 Noise robustness
In our noise simulations, $\eta\left(t\right)$ is a normally-distributed random function on a time grid with time points $t_{i}$ separated by intervals $\Delta t=5\cdot 10^{-3}/\Omega_{0}$. For simplicity, we use cubic splines to interpolate $\eta\left(t\right)$ between the $t_{i}$, and use a standard numerical integrator (SciPy’s odeint). This does not influence the physical results since $\Delta t\ll 2\pi/\Omega_{0}$. The symmetry-breaking steady state is stable up to a threshold noise amplitude $A_{*}$.
To relate the noise amplitude $A_{*}$ to a temperature, we calculate the internal energy of the oscillator at finite temperatures via $U=-\partial_{\tilde{\beta}}\log Z$:
$U$ $=-\frac{1}{2\pi\hbar}\frac{\partial}{\partial\left(1/k_{B}T\right)}\log\int dq\int dpe^{-\frac{H\left(p,q\right)}{k_{B}T}}$
$=\frac{1}{8}\left(\frac{\Omega^{4}K_{\frac{5}{4}}\left(\frac{\Omega^{4}}{8\beta k_{B}T}\right)}{\beta K_{\frac{1}{4}}\left(\frac{\Omega^{4}}{8\beta k_{B}T}\right)}-\frac{\Omega^{4}}{\beta}+2k_{B}T\right)$
$\approx k_{B}T-\frac{3\beta\left(k_{B}T\right)^{2}}{4\Omega^{4}}-\mathcal{O}\left(\beta^{2}\right).$
Assuming a phonon resonance frequency of 33\,THz, and extracting the internal energy $U$ corresponding to noise of amplitude $A_{*}$ (in absence of driving), we estimate that $A_{*}$ corresponds to a temperature of about 700700\,K.
Appendix.3 Pumping via Raman-active modes
In centrosymmetric materials, Raman- and IR-active modes are mutually exclusive. They can be, however, nonlinearly coupled to each other via the coupling Hamiltonian of Eq. (17), where $Q_{R}$ is the Raman active and $Q_{x}$ the IR active mode. The equations of motion for the coupled system can then be written down as
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}+\beta Q_{x}^{3}$ $=2AQ_{x}Q_{R}$
$+Z_{x}E_{0}\cos\left(\omega t\right)$ (A 1)
$\ddot{Q}_{R}+2\gamma\dot{Q}_{R}+\Omega_{R}^{2}Q_{R}+\beta_{R,2}Q_{R}^{2}+\beta_{R,3}Q_{R}^{3}.$
$\qquad=AQ_{x}^{2}+\frac{1}{2}\chi_{R}E_{0}^{2}\left(1+\cos\left(2\omega t\right)\right)$ (A 2)
We assume that the drive is in resonance with the Raman-active mode, such that $2\omega\approx\Omega_{R}$. Furthermore, we assume that $\Omega_{0}$ is distinct from $\Omega_{R}$, such that the driving is off-resonant with $\Omega_{0}$, and the direct, off-resonant driving of $Q_{x}$ via $Z_{x}$ is neglectable. We will furthermore be interested in instabilites of $Q_{x}$ induced by the Raman mode. At the instability onset, which we want to study here, $Q_{x}$ will be small, and we can neglect higher order terms in this variable. The simplified system of equations reads
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}-2AQ_{x}Q_{R}$ $=0$ (A 3)
$\ddot{Q}_{R}+2\gamma\dot{Q}_{R}+\Omega_{R}^{2}Q_{R}+\beta_{R,2}Q_{R}^{2}+\beta_{R,3}Q_{R}^{3}$
$\qquad=\frac{1}{2}\chi_{R}E_{0}^{2}\left(1+\cos\left(2\omega t\right)\right).$ (A 4)
Inverting the second equation, we find, neglecting higher harmonics
$Q_{R}\approx F_{R,0}\left(E_{0}\right)+F_{R,2}\left(E_{0}\right)\cos\left(2\omega t\right),$ (A 5)
where $F_{R,0}$ and $F_{R,2}$ are the amplitudes of DC and second harmonic (with respect to the driving frequency) components. Inserting this into the first equation gives
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\tilde{\Omega}_{0}\left(E_{0}\right)\left[1-h\left(E_{0}\right)\cos\left(2\omega t\right)\right]Q_{x}=0.$ (A 6)
Here $\tilde{\Omega}_{0}\left(E_{0}\right)=\Omega_{0}^{2}+F_{R,0}\left(E_{0}\right)$ is the effective, blue shifted resonance frequency of the IR active mode and $h\left(E_{0}\right)=2AF_{R,2}\left(E_{0}\right)$. This equation maps onto Eq. () of the main text. Showing, that a Raman-active mode can be used to trigger the symmetry-breaking instability of the IR active mode. The frequency of the resonant Raman-active mode $\Omega_{R}\approx 2\omega$ has to be reasonbly close to $\Omega_{0}$. However, some detuning is possible, because the effective resonance frequency $\tilde{\Omega}_{0}\left(E_{0}\right)$ experiences a blue shift. This blue shift can be exploted to adjust the effective resonance frequency by adjusting the driving amplitude $E_{0}$.
Similarly, an auxiliary IR-active mode with a frequency close to half the resonance $\Omega_{0}$ can be used, instead of the Raman-active mode. An advantage of using an auxiliary (Raman or IR-active mode) is that the mode will story the energy supplied by the drive over many cycles. Thus the threshold amplitude for the instability might be easier reach.
Figure A 2: Illustration of the instability condition of Eq. (11) for the same parameteres as in Fig. 1a).
Appendix.4 Collective instability of the x and y modes
The equations of motion (with the addition of a damping term) following form Eq. (18) read
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}+\beta Q_{x}\left(Q_{x}^{2}+Q_{y}^{2}\right)$ $=Z_{x}E_{x}\cos\omega t$
$\ddot{Q}_{y}+2\gamma\dot{Q}_{y}+\Omega_{0}^{2}Q_{y}+\beta Q_{y}\left(Q_{x}^{2}+Q_{y}^{2}\right)$ $=Z_{y}E_{y}\sin\omega t.$ (A 7)
$\ddot{Q}_{x/y,\mathrm{even}}+2\gamma\dot{Q}_{x/y,\mathrm{even}}$
$\qquad+\left[\tilde{\Omega}_{x/y}^{2}\pm\frac{\alpha}{2}\left(3F_{x/y,1}^{2}-F_{y/x,1}^{2}\right)\cos\left(2\omega t\right)\right]Q_{x/y,\mathrm{even}}$
$\qquad+\alpha F_{x/y,1}F_{y/x,1}\sin\left(2\omega t\right)Q_{y/x,\mathrm{even}}=0$ (A 8)
with $\tilde{\Omega}_{x/y}^{2}=\Omega_{0}^{2}+\alpha\left(3F_{x/y}^{2}+F_{y/x}^{2}\right)/2$, where $F_{y}$ is defined analogously to $F_{x}$ in Eq. (7). As above, we expect parametric resonances near $2\omega=\tilde{\Omega}_{x/y}$, where $\tilde{\Omega}_{x}\neq\tilde{\Omega}_{y}$ for $F_{x}\neq F_{y}$. For now, let us choose the case $2\omega=\tilde{\Omega}_{x}$, such that the instability occurs for the $Q_{x}\left(t\right)$ component.
The oscillating terms in Eqs. (A 8) couple harmonics with frequencies $2\omega$, $4\omega$, … and DC terms. For the stability analysis, we therefore choose the ansatz
$Q_{i,\mathrm{even}}$ $=a_{i,1}\sin\left(2\omega t\right)+a_{i,2}\sin\left(4\omega t\right)+b_{i,0}$
$+b_{i,1}\cos\left(2\omega t\right)+b_{i,2}\cos\left(4\omega t\right).$ (A 9)
Furthermore, we neglect $\gamma$ for the duration of this analysis. While $\gamma$ determines the instability threshold amplitudes of the electromagnetic fields [see Eq. (12)], its effects become less important for driving amplitudes above the threshold, i.e., for any driving amplitude, $\gamma$ can be always chosen small enough that our analysis is accurate. We again search for the instability window for the detuning $\Delta=2\omega-\tilde{\Omega}_{x}$, such that the mode amplitudes $a_{i,n}$ and $b_{i,n}$ grow exponentially for $\Delta_{\mathrm{min}}<\Delta<\Delta_{\mathrm{max}}$. At $\Delta=\Delta_{\mathrm{max}/\mathrm{min}}$, the amplitudes will be constant. The boundaries of the instability interval $\Delta_{\mathrm{max}/\mathrm{min}}$ are then found by inserting the ansatz (A 9) into Eqs. (A 8) and assuming that $a_{i,n}$ and $b_{i,n}$ are indeed constant (Landau_Lifshitz_Mechanics, ). After a lengthy calculation, in which we compare the coefficients of different harmonics after inserting the ansatz (A 9) into Eqs. (A 8), we find that, to fourth order in $F_{x,1}$ and $F_{y,1}$
$\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ $=F_{y,1}^{4}\left(\frac{\beta^{2}}{16\Omega_{0}^{3}}+\frac{287\beta^{4}F_{x,1}^{4}}{576\Omega_{0}^{7}}+\frac{47\beta^{3}F_{x,1}^{2}}{192\Omega_{0}^{5}}\right)$
$+F_{y,1}^{2}\left(\frac{77\beta^{3}F_{x,1}^{4}}{192\Omega_{0}^{5}}-\frac{5\beta^{2}F_{x,1}^{2}}{8\Omega_{0}^{3}}\right)+\frac{9\beta^{2}F_{x,1}^{4}}{16\Omega_{0}^{3}}.$ (A 10)
The full result is too long to be quoted here but is easily found using computer algebra. Eq. (A 10) is valid for small $F_{x}$ and $F_{y}$, i.e. for small driving amplitudes. It is interesting to study the behavior of $\Delta$ close to $F_{y,1}=F_{x,1}$, i.e. for nearly perfect circular polarization. Writing $F_{y,1}=F_{x,1}+F_{\epsilon}$, we find
$\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ $=-F_{\epsilon}^{3}\left(\frac{9\Omega_{0}}{8F_{x,1}^{3}}+\frac{\beta^{2}F_{x,1}}{\Omega_{0}^{3}}+\frac{51\beta}{16F_{x,1}\Omega_{0}}\right).$ (A 11)
Notice that for $F_{y,1}\geq F_{x,1}$ (we choose both amplitudes positive w.l.o.g.), we have $\Delta_{\mathrm{max}}\leq\Delta_{\mathrm{min}}$, which indicates that the system is stable. For $F_{y,1}>F_{x,1}$, the $Q_{x}$ and $Q_{y}$ components switch places, and the instability occurs for $2\omega=\tilde{\Omega}_{y}$. The analysis for this case is completely analogous with $F_{x,1}$ and $F_{y,1}$ , as well as $\Omega_{x}$ and $\Omega_{y}$ interchanged. We therefore conclude that, in general, the instability occurs either for the $Q_{x}$ or the $Q_{y}$ component, depending on whether $F_{x,1}$ or $F_{y,1}$ is larger. For $E_{x}=E_{y}$, resulting in $F_{x}=F_{y}$, we find $\Delta_{\mathrm{min}}=\Delta_{\mathrm{max}}$ and the instability cannot be reached. (Note2, ).